# Power Transform Revisited: Numerically Stable, and Federated

### v1

@@ 6.2 Numerical Experiments
We now test numerical stability of power transforms on four datasets: Blood (Yeh,, 2008), Cancer (Wolberg et al.,, 1995), Ecoli (Nakai,, 1996), and House (Montoya and DataCanary,, 2016). The first three were previously identified by Marchand et al., (2022) as unstable cases; House is a new addition with features exhibiting extreme skewness. We benchmark three methods: (1) ExpSearch, a derivative-based optimization method (Marchand et al.,, 2022) vs. our derivative-free approach (based on Brent’s method); (2) Linear-domain computation with our proposed log-domain method. (3) Naive one-pass variance computation vs. our pairwise computation in federated learning.
@@
Next, we observe that seven features cause ExpSearch to fail. Instead of finding $\lambda^{*}$, it returns either a boundary of the search interval or an arbitrary value. Figures 8 and 12 (in the Appendix) illustrate this. The left plot shows the true NLL curve and the optimum $\lambda^{*}$ (found by Brent’s method). ExpSearch’s $\lambda$ is far from optimal. The middle and right plots show the modified derivative used by ExpSearch versus the true derivative. Both are either unstable (returning arbitrary values) or overflow (returning boundary solutions). This explains ExpSearch’s failures.
@@
Finally, we evaluate federated NLL computation. We split each dataset into 100 clients and compare our variance aggregation method with naive one-pass variance computation. Figure 14 (in the Appendix) shows that our method produces smooth NLL curves, while the naive method introduces spikes that can disrupt optimization. This highlights the necessity of numerically stable aggregation in federated settings.
@@ 7 Discussion
Privacy is a key concern in federated learning. Approaches such as Secure Aggregation (Bonawitz et al.,, 2017), Trusted Execution Environments (TEE) (Sabt et al.,, 2015), and Secure Multiparty Computation (SMPC) (Lindell,, 2020) can ensure that only aggregated statistics are revealed to the server, not individual client data. Our method uses pairwise variance aggregation for numerical stability. To integrate with SMPC, the division operations in Equations (7) and (8) can be handled by treating the sample count $n$ as a public value, which is standard in federated learning and also adopted by Marchand et al., (2022). Preserving privacy during iterative optimization without revealing intermediate results in each round is a stronger requirement and we leave it for future work.
@@ 8 Conclusion
Numerical issues have long been a challenge in scientific computing, as mathematically equivalent expressions can yield vastly different results under finite-precision arithmetic. In this paper, we addressed the numerical instability of power transforms, a widely used technique for data normalization. We conducted a detailed analysis of the sources of instability and proposed numerically stable approachs that combines log-domain computation, reformulated expressions, and bounding strategies. We further extended power transforms to the federated learning setting, introducing a numerically stable variance aggregation method suitable for distributed data. Our empirical results demonstrate the effectiveness and robustness of our methods in both centralized and federated scenarios. We believe that our work not only makes power transforms more reliable in practice, but also provides insights that can be applied to a broader range of numerical stability problems in scientific computing.
@@
    When $\lambda=0$, $\ln x\geq 0$ for $x\geq 1$. Hence $\psi(\lambda,x)\geq 0$ for all $\lambda$ whenever $x\geq 1$. Similarly, for $0<x<1$, we have
@@
    The second order partial derivative of $\psi$ with respect to $x$ is
@@
    It’s clear that $\psi(\lambda,x)$ is continuous for $\lambda$ and $x$ except $\lambda=0$. We just need to prove it’s continuous at $\lambda=0$. By L’Hopital’s rule, we have
    $\lim_{\lambda\to 0}\frac{x^{\lambda}-1}{\lambda}=\lim_{\lambda\to 0}\frac{x^{\lambda}\ln x}{1}=\ln x$ (20)
@@
    We prove this by induction. Let $k=1$, then for $\lambda\neq 0$
    $\psi^{(1)}(\lambda,x)=\frac{x^{\lambda}\lambda\ln x-(x^{\lambda}-1)}{\lambda^{2}}=\frac{x^{\lambda}\ln x-\psi^{(0)}(\lambda,x)}{\lambda}$ (21)
    For $\lambda=0$, by L’Hopital’s rule, we have
    $\psi^{(1)}(0,x)$ $=\lim_{\lambda\to 0}\psi^{(1)}(\lambda,x)$ (22)
    $=\lim_{\lambda\to 0}\frac{x^{\lambda}\lambda\ln x-x^{\lambda}+1}{\lambda^{2}}$ (23)
    $=\lim_{\lambda\to 0}\frac{x^{\lambda}(\ln x)^{2}}{2}$ (24)
    $=(\ln x)^{2}/2$ (25)
    Assume that this hold for $k=n$ where $n\geq 1$, then for $k=n+1$ and $\lambda\neq 0$
@@
    $=\frac{x^{\lambda}(\ln x)^{n+1}-(n+1)\psi^{(n)}(\lambda,x)}{\lambda}$ (28)
    For $\lambda=0$, by L’Hopital’s rule, we have
    $\psi^{(n+1)}(0,x)$ $=\lim_{\lambda\to 0}\psi^{(n+1)}(\lambda,x)$ (29)
    $=\lim_{\lambda\to 0}\frac{x^{\lambda}(\ln x)^{n+1}-(n+1)\psi^{(n)}(\lambda,x)}{\lambda}$ (30)
    $=\lim_{\lambda\to 0}x^{\lambda}(\ln x)^{n+2}-(n+1)\psi^{(n+1)}(\lambda,x)$ (31)
    $=(\ln x)^{n+2}-(n+1)\lim_{\lambda\to 0}\psi^{(n+1)}(\lambda,x)$ (32)
    Since (29) is equal to (32), $\psi^{(n+1)}(0,x)=\lim_{\lambda\to 0}\psi^{(n+1)}(\lambda,x)=(\ln x)^{n+2}/(n+2)$. Thus, the recurrence relation holds for all $k\geq 1$ and $\lambda$.
@@ Appendix B Log-domain Computation
The computation of $\ln(x_{i}-\mu)$ requires extra care. A stable way is to rewrite the difference using the LSE trick (Haberland,, 2023):
@@ Appendix D Federated NLL Computation for Yeo-Johnson
Appendix E Supplementary Plots
Figure 9: ROC curves after applying different transforms (QDA and LR).
Figure 10: Effect of varying $\lambda$ on AUC scores (QDA and LR).

### v2

@@ 6.2 Numerical Experiments
Figure 10: Comparison of pairwise and naive variance aggregation.
We now test numerical stability of power transforms on four datasets: Blood (Yeh, 2008), Cancer (Wolberg et al., 1995), Ecoli (Nakai, 1996), and House (Montoya and DataCanary, 2016). The first three were previously identified by Marchand et al. (2022) as unstable cases; House is a new addition with features exhibiting extreme skewness. We benchmark three methods: (1) ExpSearch, a derivative-based optimization method (Marchand et al., 2022) vs. our derivative-free approach (based on Brent’s method); (2) Linear-domain computation with our proposed log-domain method; and (3) Naive one-pass variance computation vs. our pairwise computation in federated learning.
Next, we observe that seven features cause ExpSearch to fail. Instead of finding $\lambda^{*}$, it returns either a boundary of the search interval or an arbitrary value. Figures 9 and 14 (in the Appendix) illustrate this. The left plot shows the true NLL curve and the optimum $\lambda^{*}$ (found by Brent’s method). ExpSearch’s $\lambda$ is far from optimal. The middle and right plots show the modified derivative used by ExpSearch versus the true derivative. Both are either unstable (returning arbitrary values) or overflow (returning boundary solutions). This explains ExpSearch’s failures.
@@
Finally, we evaluate federated NLL computation. We split each dataset into 100 clients and compare our variance aggregation method with naive one-pass variance computation. Figures 10 and 16 (in the Appendix) show that our method produces smooth NLL curves, while the naive method introduces spikes that can disrupt optimization. This highlights the necessity of numerically stable aggregation in federated settings.
@@ 7 Discussion
Runtime performance is an important consideration for practical deployment. Our stable recipe in Section 4 modifies the NLL computation and is indeed slower than the naive implementation due to the use of log-domain variance computation, as shown in Table 4 (Appendix). However, for scientific software, numerical correctness must take priority over raw speed. Moreover, the runtime remains well below human-perceptible latency in all cases, making it acceptable for most practical applications. In federated settings, our current implementation of variance aggregation (Algorithm 1) processes clients sequentially. This can be further optimized by parallelizing the pairwise aggregation steps, thereby reducing overall latency without affecting numerical stability.
@@
Privacy is a key concern in federated learning. Approaches such as Secure Aggregation (Bonawitz et al., 2017), Trusted Execution Environments (TEE) (Sabt et al., 2015), and Secure Multiparty Computation (SMPC) (Lindell, 2020) can ensure that only aggregated statistics are revealed to the server, not individual client data. Our method uses pairwise variance aggregation for numerical stability. To integrate with SMPC, the division operations in Equations (7) and (8) can be handled by treating the sample count $n$ as a public value, which is standard in federated learning and also adopted by Marchand et al. (2022). Preserving privacy during iterative optimization without revealing intermediate results in each round is a stronger requirement and we leave it for future work.
Client sampling is a common technique in federated learning to handle offline or dropout clients. When a random subset of clients is selected each round, the NLL estimate becomes biased. However, normalizing the NLL by the subsample size recovers unbiasedness without affecting $\lambda$ optimization. Specifically, the Box-Cox NLL $(1-\lambda)\sum_{i}\ln x_{i}+\frac{n}{2}\ln\sigma^{2}_{\psi_{\text{BC}}}$ divided by $n$ yields unbiased estimates of both the log-mean term and log-variance term under uniform random sampling. Thus, the normalized NLL can be reliably estimated even with random client subsampling.
@@ 8 Conclusion
Numerical issues have long been a challenge in scientific computing, as mathematically equivalent expressions can yield vastly different results under finite-precision arithmetic. In this paper, we addressed the numerical instability of power transforms, a widely used technique for data normalization. We conducted a detailed analysis of the sources of instability and proposed numerically stable approaches that combines log-domain computation, reformulated expressions, and bounding strategies. We further extended power transforms to the federated learning setting, introducing a numerically stable variance aggregation method suitable for distributed data. Our empirical results demonstrate the effectiveness and robustness of our methods in both centralized and federated scenarios. We believe that our work not only makes power transforms more reliable in practice, but also provides insights that can be applied to a broader range of numerical stability problems in scientific computing.
Acknowledgement
We thank Matt Haberland for helpful discussions and feedback on our implementation in the GitHub pull request to SciPy. We also thank the anonymous reviewers for their valuable comments.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Acknowledgement
    When $\lambda=0$, $\ln x\geq 0$ for $x\geq 1$. Hence $\psi(\lambda,x)\geq 0$ for all $\lambda$ whenever $x\geq 1$. Similarly, for $0<x<1$, we have:
@@
    The second order partial derivative of $\psi$ with respect to $x$ is:
@@
    It’s clear that $\psi(\lambda,x)$ is continuous for $\lambda$ and $x$ except $\lambda=0$. We just need to prove it’s continuous at $\lambda=0$. By L’Hopital’s rule, we have:
    $\lim_{\lambda\to 0}\frac{x^{\lambda}-1}{\lambda}=\lim_{\lambda\to 0}\frac{x^{\lambda}\ln x}{1}=\ln x.$ (20)
@@
    We prove this by induction. Let $k=1$, then for $\lambda\neq 0$:
    $\psi^{(1)}(\lambda,x)=\frac{x^{\lambda}\lambda\ln x-(x^{\lambda}-1)}{\lambda^{2}}=\frac{x^{\lambda}\ln x-\psi^{(0)}(\lambda,x)}{\lambda}.$ (21)
    Assume that this hold for $k=n$ where $n\geq 1$, then for $k=n+1$ and $\lambda\neq 0$:
@@
    $=\frac{x^{\lambda}(\ln x)^{n+1}-(n+1)\psi^{(n)}(\lambda,x)}{\lambda}.$ (24)
    For $\lambda=0$, recall the Taylor expansion of $x^{\lambda}$ at $\lambda=0$ is:
    $x^{\lambda}=1+\lambda\ln x+\frac{(\lambda\ln x)^{2}}{2!}+\frac{(\lambda\ln x)^{3}}{3!}+\cdots=\sum_{m=0}^{\infty}\frac{(\lambda\ln x)^{m}}{m!}.$ (25)
    Thus,
    $\psi(\lambda,x)=\frac{x^{\lambda}-1}{\lambda}=\ln x+\frac{\lambda(\ln x)^{2}}{2!}+\frac{\lambda^{2}(\ln x)^{3}}{3!}+\cdots=\sum_{m=1}^{\infty}\frac{\lambda^{m-1}(\ln x)^{m}}{m!}.$ (26)
    Therefore, the $k$-th order partial derivative of $\psi$ with respect to $\lambda$ is:
    $\frac{\partial^{k}\psi(\lambda,x)}{\partial\lambda^{k}}=\sum_{m=k+1}^{\infty}\frac{(m-1)!\lambda^{m-k-1}(\ln x)^{m}}{m!}=\sum_{m=k+1}^{\infty}\frac{\lambda^{m-k-1}(\ln x)^{m}}{m}.$ (27)
    Finally, at $\lambda=0$, we have:
    $\frac{\partial^{k}\psi(\lambda,x)}{\partial\lambda^{k}}\Biggr|_{\lambda=0}=\frac{(\ln x)^{k+1}}{k+1}.$ (28)
    This completes the induction.
@@ Appendix B Log-domain Computation
The computation of $\ln(x_{i}-\mu)$ requires extra care. A stable way is to rewrite the difference using the LSE trick (Haberland, 2023):
@@ Appendix D Federated NLL Computation for Yeo-Johnson
Appendix E Supplementary Tables and Plots
Table 3: Dataset statistics.
Datasets # Row # Col
Adult 33K 14
Bank 45K 16
Credit 30K 24
Figure 11: ROC curves after applying different transforms (LR and XGB).
Figure 12: Effect of varying $\lambda$ on AUC scores (LR and XGB).
@@ Appendix E Supplementary Tables and Plots
Table 4: Runtime comparison between naive and our method for Box-Cox NLL computation.
Data size Time (Naive) Time (Ours)
[…]